OpenStudy (anonymous):

rewrite the expression without using absolute value notation. the absolute value is around the 1- sqrt2 then +1

OpenStudy (anonymous):

|1- sqrt2| + 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then its sqrt2 - 1 + 1 = sqrt 2

OpenStudy (anonymous):

do u follow?

OpenStudy (anonymous):

I have to write all the steps out so I am trying to figure out what is in between

OpenStudy (anonymous):

This is what I have so far

OpenStudy (anonymous):

\[1-\sqrt{2}<1, \left| 1-\sqrt{2} \right|=?\]

OpenStudy (anonymous):

Do I do this next \[1-\sqrt{2}=?\]

OpenStudy (anonymous):

well 1 - sqrt2 < 0 ie negative so | 1 - sqrt2| is positive and equal to sqrt2 - 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hope this helps see = it = 1 - 1.414 aprox = -0.414

OpenStudy (anonymous):

ok you lost me

OpenStudy (anonymous):

i got it now thanks